OpenStudy (anonymous):

a distribution center receives shipments of a product from three different factories in the quantaties of 60,35, and 25. three times a product is selected at randomn, each time without replacement. find the probability that (a) all three products came from the second factory and (b) none of the three products came from the second factory.

OpenStudy (campbell_st):

well isn't the probability P(2nd Factory) = 35/120 and without replacement the next selection P(2nd Factory) = 24/119 and repeat for the 3rd selection so the probability of 3 from 2nd factory will be the product of the probabilities. and not from factory 2 will be the complementary event

OpenStudy (anonymous):

no idea what ur talking about...

OpenStudy (campbell_st):

ok... then it make it hard to solve the problem

OpenStudy (campbell_st):

but the total to choose from is 60 + 35 + 25 = 120 so there are 120 products to select from... does that make sense..?

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

ok... so lets assume factory 1 = 60, factory 2 = 35 and factory 3 = 25 whats the probability of selecting a product from factory 2?

OpenStudy (anonymous):

35/120?

OpenStudy (campbell_st):

ok that correct... now the term no replacement means that there is 1 less product from factory 2 and 1 less product in the total. what would the probability of the next selection coming from factory 2 be?